Suppose your friend Bill is considering buying a Ford F150 Lariat truck. He is going to make monthly payments (at the end of each month) of $819.48 on a loan of $47,000. If the annual percentage rate is 7.80% how long in years is the loan for? Please show ALL work with finance formulas.
Answer #1

We can solve this using PV of annuity formula:

PV of annuity = P(1-((1+r)^-n))/r

r

0.078/12=0.0065

n

?

P

819.48

PV

47,000

Where r is the rate and n is the number of period. P is the payment per period

Substituting values we have:

47000= 819.48(1-((1+0.0065)^-n))/0.0065

n=72 months or 6 years